OpenStudy (anonymous):

A garden table and a bench cost $554 combined. The garden table costs $96 less than the bench. What is the cost of the bench?

OpenStudy (kidrah69):

would you just subtract? $554 - $96?

OpenStudy (anonymous):

No. incorrect

OpenStudy (kidrah69):

ok lets set variables G = garden tool B = bench

OpenStudy (kidrah69):

G+B=$554

OpenStudy (kidrah69):

so 554 = B+ (B- 96)

OpenStudy (kidrah69):

Do you get it so far?

OpenStudy (kidrah69):

554 = x + x - 96

OpenStudy (kidrah69):

im sure you can solve the rest by yourself :)